# Towards Autonomous Indoor Parking: A Globally Consistent Semantic SLAM System and A Semantic Localization Subsystem

### v1

@@ Abstract
We propose a globally consistent semantic SLAM system (GCSLAM) and a semantic-fusion localization subsystem (SF-Loc), which achieves accurate semantic mapping and robust localization in complex parking lots. Visual cameras (front-view and surround-view), IMU, and wheel encoder form the input sensor configuration of our system. The first part of our work is GCSLAM. GCSLAM introduces a novel factor graph for the optimization of poses and semantic map, which incorporates innovative error terms based on multi-sensor data and BEV (bird’s-eye view) semantic information. Additionally, GCSLAM integrates a Global Slot Management module that stores and manages parking slot observations. SF-Loc is the second part of our work, which leverages the semantic map built by GCSLAM to conduct map-based localization. SF-Loc integrates registration results and odometry poses with a novel factor graph. Our system demonstrates superior performance over existing SLAM on two real-world datasets, showing excellent capabilities in robust global localization and precise semantic mapping.
@@ I INTRODUCTION
Simultaneous Localization and Mapping (SLAM) is a fundamental task in robotics and autonomous driving. Although SLAM has been extensively researched and applied in various scenarios, its application in indoor parking environments presents unique challenges that remain unsolved. In indoor parking lots, the absence of GNSS signals makes it impossible to directly obtain high-precision vehicle poses. Moreover, low-cost sensors such as IMUs and cameras are preferred in commercial applications. These sensors face challenges in indoor parking environments, due to the unique lighting conditions, repeated structural features, and similar textures across different regions. All of these situations can hinder the effectiveness of visual SLAM.
Recently, some methods [1, 2, 3, 4] have shown capability in performing SLAM within indoor parking environment for autonomous valet parking (AVP) tasks. These methods use learning-based semantic information extracted from surround-view cameras to address the issue of challenging lighting conditions. However, their system structures lack enough constraints for optimization in the parking environment, which leads to poor performance in large and intricate parking lots. Additionally, test scenes of AVP SLAM methods are relatively simple, typically involving only a few rows of parking slots. However, real-world parking lots are complicated, characterized by a high density of parking slots and usually cover large areas that require long-distance driving within the lot. In such scenes, existing AVP SLAM methods are more prone to misalignment and distortion. Furthermore, these methods fail to handle noise and false detections in semantic segmentation, leading to inaccurate mapping results and decreased localization accuracy.
@@
Figure 2: Our system consists of three parts. The Global Slot Management module provides semantic information based on the BEV perception module. Leveraging this semantic information, GCSLAM generates three semantic terms, including Registration Error Term, Adjacent Error Term, and Global Vertical Error Term. Subsequently, the factor graph $\mathcal{G}_{\text{GCSLAM}}$ jointly optimizes the poses and semantic map to achieve a precise global map. GCSLAM is only performed when entering a new parking lot for the first time. Once the map is created, SF-Loc provides precise localization results through a novel factor graph $\mathcal{G}_{\text{SF-Loc}}$.
To address these issues, we propose a globally consistent semantic SLAM system GCSLAM. GCSLAM consists of an innovative factor graph and a global slot management module based on odometry, semantic segmentation, and slot detections. Inspired by MOFIS [3], we construct a factor graph with two types of nodes: pose and slot. To address the slot-misalignment issues, we design a new slot association method to accurately determine the relationships between parking slots. In terms of the slot distortion problem, we introduce an adjacent error term to reduce the gap between adjacent slots caused by false detection and noise. Additionally, we design a global vertical error term to constrain the orientation of parking slots. Moreover, we introduce a global slot management module to store slot observations and update them accordingly. This module includes an unstable slot filtering strategy and a slot update strategy, which can address the false detections and noise problems caused by BEV perception module.
For an unknown parking lot, once we establish a complete global map using GCSLAM, subsequent revisits do not require rerunning GCSLAM. Instead, we activate the map-based localization subsystem. Map-based localization typically offers higher accuracy and faster performance due to the advantage of prior scene knowledge. However, unlike 3D point cloud maps obtained from LiDAR, visual feature maps are sparse, which results in lower registration accuracy. In some extremely sparse regions, this can lead to convergence failure of the registration. Existing methods lack improvements for these problems. Therefore, we propose a semantic-fusion localization subsystem (SF-Loc). To address the unstable jumps in the ICP and the accumulated drift of the odometry, SF-Loc combines semantic ICP registration and odometry based on factor graph optimization to achieve robust and accurate localization.
We conduct experiments in complex parking lots with high parking slot density. The globally consistent mapping and localization results demonstrate the robustness and effectiveness of our GCSLAM. Furthermore, we test our localization subsystem SF-Loc on the global map established by GCSLAM, achieving decimeter-level global accuracy.
@@
  • We propose a globally consistent semantic SLAM system GCSLAM, which is based on factor graph optimization with innovative slot representation and novel geometry semantic combined error terms for constraints.
@@
  • We propose a map-based localization subsystem SF-Loc, which fuses semantic-ICP results and odometry constraints using factor graph optimization.
@@ II RELATED WORK
Nevertheless, these methods are unable to perform SLAM within indoor parking slots for AVP tasks due to the complex conditions of indoor environments, such as limited distinctive features and complex lighting conditions. To solve these problems, some works [1, 2, 3, 4] are proposed. These works all utilize bird’s-eye view (BEV) images as input, which can provide rich ground features to tackle the issue of limited distinctive features in parking lots. AVP-SLAM [1] uses semantic segmentation to annotate parking spaces, ground markings, speed bumps, and other information in the images, as the segmentation method can effectively adapt to complex lighting conditions. This semantic information is added to a global map, which is then used for registration-assisted localization. However, their map is a pure point cloud map used for registration, without recording each parking lot independently and lacking important attribute information such as the location and angle of each parking slot. Zhao et al. [4] utilize parking slot detector [17] to detect the entry points of parking slots and combine the observations of parking slots with odometry to construct new positioning factors. However, this method does not maintain an overall parking space map. Instead, it primarily uses the map as an auxiliary tool for localization. VISSLAM [2] adds constraints between parking spaces, combining odometry information to propose a parking slots management algorithm that improves mapping results. Subsequent work MOFISSLAM [3] incorporates sliding window optimization, achieving higher localization accuracy and improved mapping results.
Nevertheless, existing methods are sensitive to noise and perform poorly in complex parking lots. To address this, we propose a novel factor graph for indoor parking SLAM, improving both robustness and accuracy.
@@ III METHOD
Our system adopts multiple sensors as input, including a front-view camera, IMU, wheel encoder, and four surround-view cameras. The overall framework of our work is shown in Fig. 2. The first part of our work is the SLAM system GCSLAM. GCSLAM integrates three modules: global slot management module, odometry, and factor graph optimization. The odometry module is loosely coupled with other modules, making it replaceable and enhancing the system’s flexibility and usability. In this paper, we employ VIW [18] as our odometry module. The global slot management module includes BEV perception module and slot association. Our BEV perception module is a multi-task framework based on [19, 20, 21]. It takes BEV images as input to generate semantic segmentation results (ground markings) and slot detection results (parking boundary endpoints) in real-time, using a unified backbone network with different output heads for each specific task. Moreover, this global slot management module registers detection results to global slots and performs slot association. Based on the odometry poses, semantic information, and slot association results, factor graph optimization can achieve precise pose estimation and global semantic map construction. After establishing a global semantic map, the second part of our work, localization subsystem SF-Loc, fuses odometry poses with semantic registration results for map-based localization.
III-A Factor Graph with Semantic Slot Node
We view the SLAM task as a factor graph optimization problem, aiming to estimate the accurate pose of keyframes. The keyframe is selected based on the inter-frame distance provided by the odometry. A factor graph consists of nodes and edges, where nodes represent the variables to be optimized, and edges are error terms that constrain the nodes. As depicted in Fig. 3, GCSLAM constructs the factor graph $\mathcal{G}_{\text{GCSLAM}}$ using two types of nodes $\{N_{pose}^{i}\},\{N_{slot}^{k}\}$ and four types of edges $\{E_{odo}^{i,j}\},\{E_{reg}^{k,i}\},\{E_{adj}^{k,p}\},\{E_{vert}^{k,p}\}$:
$\small\mathcal{G}_{\text{GCSLAM}}=\{\{N_{pose}^{i}\},\{N_{slot}^{k}\},\{E_{odo}^{i,j}\},\{E_{reg}^{k,i}\},\{E_{adj}^{k,p}\},\{E_{vert}^{k,p}\}\}.\mathrel{\normalsize}$ (1)
@@ III-A1 Pose Node
The pose node $N_{pose}^{i}$ stores 3 degree-of-freedom (DoF) car pose ($x,y,\theta$) of frame $i$ in world coordinates since our SLAM system assumes a planar parking lot. We initialize the pose node $N_{pose}^{i}$ using estimated pose $T_{odo}^{i}$ provided by the odometry module. This module runs as a separate thread.
@@ III-A2 Slot Node
$T^{i}=N_{pose}^{i-1}({T_{odo}^{i-1}})^{-1}{T_{odo}^{i}}.$ (2)
Then, we perform slot association to determine the global ID $k$ of this observed slot. This global slot is denoted as $S_{k}$. We record the observation of $S_{k}$ in the current frame $i$ as $O_{k}^{i}$, representing the midpoint coordinate of $S_{k}$’s entry edge in the car coordinate of frame $i$. The 3DoF pose of $S_{k}$’s entry edge midpoint in the world coordinate is denoted as the slot node $N_{slot}^{k}$.
Figure 3: The factor graph structure of our SLAM system. The factor graph consists of two types of nodes: Pose Nodes and Slot Nodes. Four types of error terms are used to constrain the nodes. The prior factor is provided by odometry and only applies to the first pose node, determining the absolute pose of the first pose.
@@ III-A3 Odometry Error Term (OET)
We construct OET between $N_{pose}^{i}$ based on the odometry module. The specific form for OET is:
$E_{odo}^{i,j}=({N_{pose}^{i}})^{-1}({N_{pose}^{i})^{-1}}-({T_{odo}^{i}})^{-1}T_{odo}^{i-1}.$ (3)
@@ III-A4 Registration Error Term (RET)
RET constrains the relationship between $N_{slot}^{k}$ and $N_{pose}^{i}$. By transforming the observation $O_{k}^{i}$ to the world coordinate through $T_{odo}^{i}$ and comparing it with $N_{slot}^{k}$, we can establish RET as follows:
$E_{reg}^{k,i}={T^{i}}{O_{k}^{i}}-{N_{slot}^{k}}.$ (4)
@@ III-A5 Adjacent Error Term (AET)
When the $i$-th keyframe arrives, all $O_{k}^{i}$ are traversed. If the distance $d_{k,p}=||O_{k}^{i}-O_{p}^{i}||$ between two slots $S_{k}$ and $S_{p}$ is less than a specified threshold (2.5m), they are considered adjacent. AET $E_{adj}^{k,p}$ is established between adjacent slots. We use this error term to ensure that the directions of adjacent slots are consistent and there are no gaps between them. The specific form for AET is:
$E_{adj}^{k,p}=\frac{\overrightarrow{w_{k}}+\overrightarrow{w_{p}}}{2}-[(N_{slot}^{k})_{1:2}-(N_{slot}^{p})_{1:2}],$ (5)
where $\overrightarrow{w_{k}}$ represents the entry edge of parking slot $S_{k}$, as shown in Fig. 4(d). $(N_{slot}^{k})_{1:2}$ denotes the first two dimensions of $N_{slot}^{k}$, representing the $xy$ coordinates of the slot.
@@ III-A6 Global Vertical Error Term (GVET)
$\begin{split}E_{vert}^{k,p}=min\{&|[(N_{slot}^{k})_{1:2}-(N_{slot}^{p})_{1:2}]\cdot\overrightarrow{d_{global}}|,\\ &|[(N_{slot}^{k})_{1:2}-(N_{slot}^{p})_{1:2}]\cdot\overrightarrow{d_{global}}^{\perp}|\}.\end{split}$ (6)
@@ III-B1 Slot Association
Once a global slot $S_{k}$ is associated with $S_{obs}$, we increment the count of observation frames $c_{O}^{k}$ for this slot. By recording the observation frequency of each global slot, we can exclude low-frequency slots as false detections. This filter strategy can effectively mitigate the noise from the BEV perception module. The specific filter logic is illustrated in Alg. 1. When a new keyframe arrives, we perform slot association to all observed slots. Then, we iterate through all unstable global slots $S_{k}$. First, we increase $c_{E}^{k}$ (the frame count since $S_{k}$ is created,). If $S_{k}$ is observed in current frame, we increase $c_{O}^{k}$ (the frame count in which $S_{k}$ is observed). If $c_{O}^{k}>9$ we label $S_{k}$ as stable; otherwise, we continue to the next unstable slot. If an $S_{k}$ has $c_{E}^{k}>30$ and is still not labeled as stable, we consider it a false detection and delete the slot.
@@ III-B2 Slot Updating
We update the global slot and its weight by:
$\begin{split}&S_{k}^{\prime}=\frac{S_{k}\times w_{k}\times(c_{O}^{k}-1)+S_{obs}}{w_{k}\times(c_{O}^{k}-1)+w_{k}^{i}},\\ &w_{k}^{\prime}=\frac{w_{k}\times(c_{O}^{k}-1)+w_{k}^{i}}{c_{O}^{k}}.\end{split}$ (9)

### v2

@@ Abstract
We propose a globally consistent semantic SLAM system (GCSLAM) and a semantic-fusion localization subsystem (SF-Loc), which achieves accurate semantic mapping and robust localization in complex parking lots. Visual cameras (front-view and surround-view), IMU, and wheel encoder form the input sensor configuration of our system. The first part of our work is GCSLAM. GCSLAM introduces a semantic-constrained factor graph for the optimization of poses and semantic map, which incorporates innovative error terms based on multi-sensor data and BEV (bird’s-eye view) semantic information. Additionally, GCSLAM integrates a Global Slot Management module that stores and manages parking slot observations. SF-Loc is the second part of our work, which leverages the semantic map built by GCSLAM to conduct map-based localization. SF-Loc integrates registration results and odometry poses with a novel factor graph. Our system demonstrates superior performance over existing SLAM on two real-world datasets, showing excellent capabilities in robust global localization and precise semantic mapping.
@@ I Introduction
Simultaneous Localization and Mapping (SLAM) is a fundamental task in robotics and autonomous driving. Although SLAM has been extensively researched and applied in various scenarios, its application in indoor parking environments presents unique challenges that remain unsolved. In indoor parking lots, the absence of GNSS signals makes it impossible to directly obtain high-precision vehicle poses. Moreover, low-cost sensors such as IMUs and cameras are preferred in commercial applications. These sensors face challenges in indoor parking environments, due to the repeated structural features and similar textures across different regions, which hinder the effectiveness of visual SLAM.
Recently, some methods [1, 2, 3, 4] have shown capability in performing SLAM within indoor parking environments for autonomous valet parking (AVP) tasks. However, these SLAM systems fail to establish adequate constraints for optimization in environments with highly repetitive structure, which leads to poor performance in large and intricate parking lots. Additionally, test scenes of these SLAM methods are relatively simple, typically involving only a few rows of parking slots. However, real-world parking lots are complicated, characterized by a high density of parking slots and usually cover large areas that require long-distance driving within the lot. In such scenes, existing methods suffer from accumulative drift in tracking, leading to misalignment and distortion in mapping. Furthermore, these methods fail to handle noise and false detections in semantic segmentation, leading to inaccurate mapping results and decreased localization accuracy.
@@
Figure 2: Our system consists of three parts. The Global Slot Management module provides semantic information based on the BEV perception module. By leveraging this semantic information, GCSLAM generates three semantic terms, including Registration Error Term, Adjacent Error Term, and Global Vertical Error Term. Subsequently, the semantic-constrained factor graph $\mathcal{G}_{\text{GCSLAM}}$ jointly optimizes the poses and semantic map to achieve a precise global map. GCSLAM is only performed when entering a new parking lot for the first time. Once the map is created, SF-Loc provides precise localization results through a novel factor graph $\mathcal{G}_{\text{SF-Loc}}$.
To address these issues, we propose a globally consistent semantic SLAM system GCSLAM, which consists of a semantic-constrained factor graph and a global slot management module designed for semantic mapping and robust tracking in parking lots. Specifically, we introduce a new slot association strategy to accurately determine the relationships between parking slots. Moreover, we design an adjacent error term and a global vertical error term in our semantic-constrained factor graph to mitigate the slot misalignment and distortion problem. The former error term is capable of reducing the gap between adjacent slots and the latter constrains parking slots to maintain correct orientations, which are caused by accumulative drift and noise detections. Furthermore, we introduce a global slot management module to store slot observations and update them accordingly. This module includes an unstable slot filter and a slot updater, which can address the false detections and noise problems caused by BEV perception module.
Apart from GCSLAM, we propose a semantic-fusion localization subsystem (SF-Loc) for map-based localization. For an unknown parking lot, after constructing a complete global map using GCSLAM, we directly activate SF-Loc for map-based localization during subsequent revisits. Such subsystem typically provide higher accuracy and faster performance as it leverages prior scene knowledge established by GCSLAM. In SF-Loc, we design a novel factor graph that integrates semantic ICP registration and odometry for optimization. This design enables robust tracking in challenging scenarios, particularly addressing convergence failures in feature-sparse regions.
@@
  • We propose a globally consistent semantic SLAM system GCSLAM, which is based on a semantic-constrained factor graph with novel semantic error terms for constraints, as well as innovative slot representation for joint optimization.
@@
  • We propose a map-based localization subsystem SF-Loc, which integrates semantic-ICP registration and odometry constraints using factor graph optimization.
@@ II RELATED WORK
However, these methods are unable to perform SLAM within indoor parking slots for AVP tasks due to the complex conditions of indoor environments, such as limited distinctive features and complex lighting conditions. To solve these problems, some works [1, 2, 3, 4] are proposed. These works all utilize bird’s-eye view (BEV) images as input, which can provide rich ground features to tackle the issue of limited distinctive features in parking lots. AVP-SLAM [1] uses semantic segmentation to annotate parking spaces, ground markings, speed bumps, and other information in the images, as the segmentation method can effectively adapt to complex lighting conditions. This semantic information is added to a global map, which is then used for registration-assisted localization. However, their map is a pure point cloud map used for registration, without recording each parking lot independently and lacking important attribute information such as the location and angle of each parking slot. Zhao et al. [4] utilize parking slot detector [18] to detect the entry points of parking slots and combine the observations of parking slots with odometry to construct new positioning factors. However, this method does not maintain an overall parking space map. Instead, it primarily uses the map as an auxiliary tool for localization. VISSLAM [2] adds constraints between parking spaces, combining odometry information to propose a parking slots management algorithm that improves mapping results. Subsequent work MOFISSLAM [3] incorporates sliding window optimization, achieving higher localization accuracy and improved mapping results.
Nevertheless, existing methods are sensitive to noise and perform poorly in complex parking lots. To address this, we propose a semantic-constrained factor graph for indoor parking SLAM, improving both robustness and accuracy.
@@ III METHOD
Our system adopts multiple sensors as input, including a front-view camera, IMU, wheel encoder, and four surround-view cameras. The overall framework of our work is shown in Fig. 2. The first part of our work is the SLAM system GCSLAM. GCSLAM integrates three modules: odometry, global slot management, and semantic-constrained factor graph optimization. The odometry module is loosely coupled with other modules, making it replaceable and enhancing the system’s flexibility and usability. In this paper, we employ VIW [19] as our odometry. The global slot management module includes BEV perception module and slot association. Our BEV perception module is a multi-task framework based on [20, 21, 22]. It takes BEV images as input to generate semantic segmentation results (ground markings) and slot detection results (parking boundary endpoints) in real-time, using a unified backbone network with different output heads for each specific task. Subsequently, the detection results are registered to global slots, followed by slot association process. Moreover, based on the odometry poses, semantic information, and slot association results, our semantic-constrained factor graph jointly optimizes pose and global semantic map to achieve robust tracking and accurate semantic mapping. After establishing a global semantic map, the second part of our work, localization subsystem SF-Loc, fuses odometry poses with semantic registration results for map-based localization.
III-A Semantic-constrained Factor Graph Optimization
We view the SLAM task as a factor graph optimization problem, aiming to estimate the accurate pose of keyframes. The keyframe is selected based on the inter-frame distance provided by the odometry. A factor graph consists of nodes and edges, where nodes represent the variables to be optimized, and edges are error terms that constrain the nodes. As depicted in Fig. 3, GCSLAM constructs the semantic-constrained factor graph $\mathcal{G}_{\text{GCSLAM}}$ using two types of nodes $\{T_{node}^{i}\},\{S^{k}\}$ and four types of edges $\{E_{odo}^{i,j}\},\{E_{reg}^{k,i}\},\{E_{adj}^{k,p}\},\{E_{vert}^{k,p}\}$:
$\small\mathcal{G}_{\text{GCSLAM}}=\{\{T_{node}^{i}\},\{S^{k}\},\{E_{odo}^{i,j}\},\{E_{reg}^{k,i}\},\{E_{adj}^{k,p}\},\{E_{vert}^{k,p}\}\}.\mathrel{\normalsize}$ (1)
@@ III-A1 Pose Node
The pose node $T_{node}^{i}$ stores 3 degree-of-freedom (DoF) car pose ($x,y,\theta$) of frame $i$ in world coordinates since our SLAM system assumes a planar parking lot. We initialize the pose node $T_{node}^{i}$ using estimated pose $T_{odo}^{i}$ provided by the odometry module. This module runs as a separate thread.
@@ III-A2 Slot Node
$T^{i}=T_{node}^{i-1}({T_{odo}^{i-1}})^{-1}{T_{odo}^{i}}.$ (2)
Then, we perform slot association to determine the global ID $k$ of this observed slot. This global slot is denoted as $S_{k}$. We record the observation of $S_{k}$ in the current frame $i$ as $O_{k}^{i}$, representing the midpoint coordinate of $S_{k}$’s entry edge in the car coordinate of frame $i$. The 3DoF pose of $S_{k}$’s entry edge midpoint in the world coordinate is denoted as the slot node.
Figure 3: The semantic-constrained factor graph structure of our SLAM system. The graph consists of two types of nodes: Pose Nodes and Slot Nodes. Four types of error terms are used to constrain the nodes. The prior factor is provided by odometry and only applies to the first pose node, determining the absolute pose of the first pose.
@@ III-A3 Odometry Error Term (OET)
We construct OET between $T_{node}^{i}$ and $T_{node}^{j}$ based on the odometry module. The specific form for OET is:
$E_{odo}^{i,j}=({T_{node}^{i}})^{-1}({T_{node}^{j})^{-1}}-({T_{odo}^{i}})^{-1}T_{odo}^{j}.$ (3)
@@ III-A4 Registration Error Term (RET)
RET constrains the relationship between $S^{k}$ and $T_{node}^{i}$. By transforming the observation $O_{k}^{i}$ to the world coordinate through $T_{odo}^{i}$ and comparing it with $S^{k}$, we can establish RET as follows:
$E_{reg}^{k,i}={T^{i}}{O_{k}^{i}}-{S^{k}}.$ (4)
@@ III-A5 Adjacent Error Term (AET)
When the $i$-th keyframe arrives, all $O_{k}^{i}$ are traversed. If the distance $d_{k,p}=||O_{k}^{i}-O_{p}^{i}||$ between two slots $S_{k}$ and $S_{p}$ is less than a specified threshold (2.5m), they are considered adjacent. AET $E_{adj}^{k,p}$ is established between adjacent slots. We use this error term to ensure that the directions of adjacent slots are consistent and there are no gaps between them. Specific form for AET is:
$E_{adj}^{k,p}=\frac{\overrightarrow{w_{k}}+\overrightarrow{w_{p}}}{2}-[(S^{k})_{1:2}-(S^{p})_{1:2}],$ (5)
where $\overrightarrow{w_{k}}$ represents the entry edge of parking slot $S_{k}$, as shown in Fig. 4(d). $(S^{k})_{1:2}$ denotes the first two dimensions of $S^{k}$, representing the $xy$ coordinates of the slot.
@@ III-A6 Global Vertical Error Term (GVET)
$\begin{split}E_{vert}^{k,p}=min\{&|[(S^{k})_{1:2}-(S^{p})_{1:2}]\cdot\overrightarrow{d_{global}}|,\\ &|[(S^{k})_{1:2}-(S^{p})_{1:2}]\cdot\overrightarrow{d_{global}}^{\perp}|\}.\end{split}$ (6)
@@ III-B1 Slot Association
Once a global slot $S_{k}$ is associated with $S_{obs}$, we increment the count of observation frames $c_{O}^{k}$ for this slot. By recording the observation frequency of each global slot, we can exclude low-frequency slots as false detections. This filter strategy can effectively mitigate the noise from the BEV perception module. The specific filter logic is illustrated in Alg. 1. When a new keyframe arrives, we perform slot association to all observed slots. Then, we iterate through all unstable global slots $S_{k}$. First, we increase $c_{E}^{k}$ (the frame count since $S_{k}$ is created,). If $S_{k}$ is observed in current frame, we increase $c_{O}^{k}$ (the frame count in which $S_{k}$ is observed). If $c_{O}^{k}>9$ we label $S_{k}$ as stable; otherwise, we continue to the next unstable slot. If an $S_{k}$ has $c_{E}^{k}>30$ and is still not labeled as stable, we consider it a false detection and delete the slot. These thresholds are selected based on comprehensive experimental data.
@@ III-B2 Slot Updating
The weighting factors in the equation were determined experimentally. We update the global slot and its weight by:
$\begin{split}&S_{k}^{\prime}=\frac{S_{k}\times w_{k}\times(c_{O}^{k}-1)+S_{obs}\times w_{k}^{i}}{w_{k}\times(c_{O}^{k}-1)+w_{k}^{i}},\\ &w_{k}^{\prime}=\frac{w_{k}\times(c_{O}^{k}-1)+w_{k}^{i}}{c_{O}^{k}}.\end{split}$ (9)